OpenStudy (anonymous):

Can someone help me simplfy 121×11^9 in exponential form plz?

OpenStudy (madhu.mukherjee.946):

121*11^9=11^2*11^9=11^9+2=11^11 :)

OpenStudy (anonymous):

Thanks @madhu.mukherjee.946 :)

OpenStudy (madhu.mukherjee.946):

your welcome darling @halia